# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 11 Mar 2026 v1

@@ 7 Implications for PRIM-LWE
$c=\inf_{m\geq 1}c_{m}(p).$
Since
$c_{m}(p)=\frac{\varphi(p-1)}{p-1}\prod_{j=1}^{m}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
is decreasing in $m$, it follows that
$\inf_{m\geq 1}c_{m}(p)=\lim_{m\to\infty}c_{m}(p)=c(p).$
Thus $c(p)$ is the dimension-uniform sampling term entering the reduction. The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
@@
In fixed dimension $n$, the exact expected number of trials is $1/c_{n}(p)\leq 1/c(p)$. By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{m\geq 1}c_{m}(p)=c(p)$ is the dimension-uniform sampling constant. Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
However, this is a modulus-dependent nuance rather than a structural obstruction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. Moreover, the explicit bounds of Section 6 show that standard choices of LWE moduli yield $c(q)$ concretely bounded away from zero. For safe primes $q=2r+1$ with $r$ an odd prime (i.e. $q\geq 7$), one has
$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right).$
For a safe prime $q=2r+1$, the prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$. Hence $c(q)$ is close to $1/2$ for large safe primes. Note, however, that the conditional safe-prime and Fermat-prime observations are now superseded by Corollary 3, which gives $\limsup c(p)=1/2$ unconditionally. On the other hand, if
$P(t):=\prod_{j\geq 1}\Bigl(1-t^{-j}\Bigr)\qquad(t>1),$
then $P(t)$ is increasing in $t$, so for every safe prime $q\geq 7$ we have $P(q)\geq P(3)=P_{3}$. Therefore
$c(q)\geq P_{3}\,\frac{r-1}{2r}.$
@@
We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility (often $q\equiv 1\pmod{2^{t}}$ for moderate $t$). Mathematically, $q\equiv 1\pmod{2^{t}}$ controls only the 2-adic part of q1q-1; it does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$. What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization. The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$. In practice, many NTT primes of cryptographic interest additionally have few small odd prime factors dividing q1q-1, and it is this latter property—not NTT-friendliness per se—that yields the good lower bounds on $c(q)$ recorded in Table 1 and Remark 7. Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
Affine determinant programs (ADPs) [26] provide a natural external determinant-based connection to prim-lwe. In the ADP obfuscation framework, a Kilian-style randomization step left- and right-multiplies by invertible matrices $R,S$ with $\det(R)\det(S)=1$ to preserve the determinant polynomial. Here the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$, not all of $M_{n}(\mathbb{F}_{q})$. Since the determinant map
@@
so rejection sampling inside $\mathrm{GL}_{n}(\mathbb{F}_{q})$ succeeds in fewer than $6.4$ trials on average. If instead one samples from all matrices $M_{n}(\mathbb{F}_{q})$, then the exact finite-$n$ primitive-root density is $c_{n}(q)$, whose limit as $n\to\infty$ is $c(q)$. We do not know whether such a restriction improves security. We only note, for context, that the ADP framework of [26] includes left- and right-randomization by invertible matrices $R,S$ satisfying $\det(R)\det(S)=1$, while the cryptanalysis of [27] targets a weakness in one of the candidate randomization steps.
@@ 8 Concluding Remarks
The proof is short and unconditional, requiring only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ product formula; the primorial prime hypothesis is entirely unnecessary. We have further established several results that illuminate the finer structure of the density $c(p)$:
@@
    The quantity $c(p)$ possesses a continuous purely singular limiting distribution over the primes, this law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, and its support is exactly $0,\tfrac{1}{2}]$. Equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$.
@@
    For many practical moduli of cryptographic interest—including the current NIST-standardized moduli and other NTT-friendly primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$. Accordingly, the prim-lwe reduction has a finite, explicitly computable constant-factor sampling term for every fixed prime modulus, and this term is small in the concrete parameter regimes treated here.
On the cryptographic side, our results show that the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but this represents a quantitative sensitivity to modulus selection, not a qualitative failure of the reduction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. The primes typically chosen for lattice-based cryptography are NTT-friendly, meaning $q\equiv 1\pmod{2^{t}}$ for moderate $t$. While NTT-friendliness alone does not imply a good bound on $c(q)$, the NTT condition is harmless with respect to the 2-part of q1q-1 (the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$), and many practical NTT primes additionally have few small odd prime factors dividing q1q-1; it is this controlled factorization structure that yields the good concrete bounds on $c(q)$. For the two NIST-standardized moduli $q=3329$ (ML-KEM) and $q=8380417$ (ML-DSA), the dimension-uniform expected rejection-sampling overheads $1/c(q)$ are at most $2.17$ and $3.42$, respectively; in fixed dimension $n$, the exact expected overheads are $1/c_{n}(q)\leq 1/c(q)$. As a simple conservative explicit estimate, for any prime $q>2^{30}$ one has $1/c(q)\leq 1.79\log q$. The sharper universal pointwise bound is Corollary 7, which gives $1/c(q)=O(\log\log q)$; equivalently, the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$. Thus our results do not by themselves indicate a practical weakness. They quantify how modulus choice affects the reduction loss in the prim-lwe framework.
@@ Appendix A Positive Density Near $1/2$
Step 3: Mean-value estimate for $S(p)$. Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
@@
Since $q>R$, we have $q\nmid M_{R}$. The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is one, and modulo $q$ the residue class is forced to be $1$. Because $\varphi(M_{R}q)=\varphi(M_{R})(q-1)$, we obtain from Theorem 5 that
@@
uniformly for $R<q\leq Q$. (Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.)

### 18 Mar 2026 v2

@@ 7 Implications for PRIM-LWE
$c=\inf_{n\geq 1}c_{n}(p).$
As shown in Section 2, $c_{n}(p)$ is decreasing in $n$ and $\inf_{n\geq 1}c_{n}(p)=c(p)$, so $c(p)$ is the dimension-uniform sampling term entering the reduction. The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
@@
In fixed dimension $n$, the exact expected number of trials is $1/c_{n}(p)\leq 1/c(p)$. By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{n\geq 1}c_{n}(p)=c(p)$ is the dimension-uniform sampling constant. Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
However, this is a modulus-dependent nuance rather than a structural obstruction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. Moreover, the explicit bounds of Section 6 show that the specific moduli surveyed in Table 2—which cover the NIST standards and several widely deployed FHE primes—all have $c(q)$ concretely bounded away from zero.
Corollary 3 gives $\limsup c(p)=1/2$ unconditionally, so there is no shortage of primes for which $c(p)$ is close to its maximal value. As a concrete illustration, for safe primes $q=2r+1$ with $r$ an odd prime and $q\geq 7$, one has
$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right)\geq P_{3}\,\frac{r-1}{2r}.$
The prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$, so $c(q)$ approaches $1/2$ for large safe primes.
@@
We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility. Concretely, this typically means $q\equiv 1\pmod{2^{t}}$ for moderate $t$. Mathematically, this condition controls only the 2-adic part of q1q-1. It does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$. What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization. The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$. Among the implementation moduli surveyed in Table 2, every prime satisfies $\omega(q-1)\leq 6$. It is this controlled factorization that yields the good lower bounds on $c(q)$ recorded in Remark 5, and Tables 1 and 2. Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
Affine determinant programs (ADPs) [36] provide a natural external determinant-based connection to prim-lwe. In the ADP obfuscation framework, a Kilian-style randomization step left- and right-multiplies by invertible matrices $R,S$ with $\det(R)\det(S)=1$ to preserve the determinant polynomial. Because the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$ rather than all of $M_{n}(\mathbb{F}_{q})$, the relevant primitive-root fraction reduces to $\varphi(q-1)/(q-1)$: the matrix-count Euler product $\prod_{j=1}^{n}(1-q^{-j})$ is already absorbed into the $\mathrm{GL}_{n}$ normalization. Since the determinant map
@@
so rejection sampling inside $\mathrm{GL}_{n}(\mathbb{F}_{q})$ succeeds in fewer than $6.4$ trials on average. If instead one samples from all matrices $M_{n}(\mathbb{F}_{q})$, then the exact finite-$n$ primitive-root density is $c_{n}(q)$, whose limit as $n\to\infty$ is $c(q)$. While our results do not establish a security advantage for imposing primitive-root determinant constraints, they confirm that the associated sampling overhead is mild whenever q1q-1 has controlled factorization.
@@ 8 Concluding Remarks
The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula. We have further established several results that illuminate the finer structure of the density $c(p)$:
@@
    The quantity $c(p)$ has a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$. This full-support conclusion appears to be new. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, as shown by a transport lemma, and is purely singular. Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p. 263] noted that it carries over to shifted primes.
@@
    For the implementation moduli surveyed in Table 2—including the current NIST-standardized moduli and several widely used FHE primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$. Accordingly, the prim-lwe reduction has a finite, explicitly computable constant-factor sampling term for every fixed prime modulus, and this term is small in the concrete parameter regimes treated here.
On the cryptographic side, the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but as discussed in Section 7, this represents a quantitative sensitivity to modulus selection rather than a structural obstruction. Table 2 confirms that every implementation modulus we examined has $\omega(q-1)\leq 6$ and $1/c(q)\leq 4.00$, and the sharper universal pointwise bound of Corollary 7 gives $1/c(q)=O(\log\log q)$.
@@ Appendix A Positive Density Near $1/2$
Step 3: Mean-value estimate for $S(p)$. Throughout Steps 3 and 4, the parameter $R$ (and hence $\varepsilon$) is fixed; all limits are as $x\to\infty$. Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
@@
Since $q>R$, we have $q\nmid M_{R}$. The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is exactly a single admissible class, and modulo $q$ the residue class is forced to be $1$. Because $\varphi(M_{R}q)=\varphi(M_{R})(q-1)$, we obtain from Theorem 5 that
@@
uniformly for $R<q\leq Q$. Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.